# Residual Random Neural Networks

### v2

@@ Abstract
The single-layer feedforward neural network with random weights is a recurring motif in the neural networks literature. The advantage of these networks is their simplified training, which reduces to solving a ridge-regression problem. However, a general assumption is that these networks require a large number of hidden neurons relative to the dimensionality of the data samples, in order to achieve good classification accuracy. Contrary to this assumption, here we show that one can obtain good classification results even if the number of hidden neurons has the same order of magnitude as the dimensionality of the data samples, if this dimensionality is reasonably high. We also develop an efficient iterative residual training method for such random neural networks, which significantly improves their classification accuracy. Moreover, we also describe an encryption (obfuscation) method which can be used to protect both the data and the neural network model.
@@ 1 Introduction
Randomization of the weights in the single-layer feedforward neural network (SLFNN) is a recurring motif in the neural networks research. As a consequence, the "benefits of randomization" have been discovered (and rediscovered) many times by different authors in various contexts [1]-[9]. The main idea is the fixed random initialization of the weights in the hidden layer of SLFNN, such that the training is reduced to computing the ridge-regression solution of the linear system corresponding to the hidden layer outputs and the classification targets. Therefore, the key advantage of these random neural networks is their simplified training algorithm. Also, it has been shown that despite of the randomness of their hidden weights layer, these simple networks are still capable of a good approximation of any non-constant piecewise continuous function [5].
A general assumption is that these random-SLFNNs require a large number of hidden neurons $J$ relative to the dimensionality of the samples $M$, in order to achieve good classification accuracy. Typically, for higher accuracy applications the number of hidden neurons is required to be an order of magnitude larger (or even more) than the dimensionality of the data samples ($J\gg M$) [1]-[9]. Thus, the classification accuracy of random-SLFNNs depends on two parameters, the dimensionality of the space $M$ and the number of the hidden random neurons $J$. Therefore, the application of random-SLFNNs to very-high-dimensional data becomes problematic, since by increasing the dimensionality of the data $M$, the number of hidden neurons $J$ is also expected to increase by an order of magnitude, or even more.
@@
Using the above intuition, here we derive a simple strategy to improve the classification accuracy of random-SLFNNs by increasing the dimensionality of the data samples $M$. This approach results in a reduced number of required hidden neurons $J$, such that both $M$ and $J$ have the same order of magnitude. We further exploit this idea by developing an efficient and low complexity iterative method, that can be used to train a residual random neural network (RRNN) model, significantly improving the classification accuracy. Moreover, we also describe an encryption (obfuscation) method based on orthonormal random matrices, which can be used to protect both the data and the neural network model.
@@ 4.1 Data
such that all the vectors involved in the computation lie on the $M-$dimensional unit hypersphere $\mathcal{S}^{M-1}(1)$. An optional initial step which may improve the classification is also $x_{n}\leftarrow\sqrt{x_{n}}$ [16]. Thus, the random projection of a data sample is simply the cosine of the angle between the sample and the random projection vector. Also, the non-linear activation function used in all the numerical experiments is the hyperbolic tangent. Therefore, after accounting for the above normalization procedure, the matrices $H$ and $\tilde{H}$ from random-SLFNN equations can be rewritten as following:
@@ 5.1 Additional information
Following the conclusion of the preliminary numerical experiments, we hypothesize that by adding more information about the data samples, such that we can further increase the size of each sample, then we can expect an increase in the classification accuracy. A simple way to add more information is to compute some non-linear function of the data sample, and appended it to the initial data sample. Obviously, such a function must be non-linear, since a linear function will make the data matrix singular. For example, a simple approach is to add the absolute value of the fast Fourier transform (FFT) [16]
@@ 5.2 Random-SLFNN with additional information
Let us now repeat the previously described experiment, by adding the additional FFT information. The results shown in Table 3 are surprisingly good, reaching an accuracy of $\eta=98.63\%$ for MNIST, and respectively $\eta=90.13\%$ for fMNIST, when $s=3$ and $q=2$. In all the considered cases the regularization parameter was set to $\lambda=0$ for MNIST, and $\lambda=10$ for fMNIST.
@@ 6.1 The Algorithm
6.2 Numerical results
In Figures 1 and 2 we show the numerical results obtained using the RRNN approach for the MNIST, and respectively the fMNIST datasets. One can see that for $s=2$, $q=2$ and $T^{*}=15$ iteration steps (or layers), the classification accuracy reaches $\eta\simeq 99.12\%$ for MNIST ($\lambda=0$, $\mu=1/2$), and respectively $\eta\simeq 91.41\%$ for fMNIST ($\lambda=10$, $\mu=1/2$). These results are surprisingly good for such a simple method, which as shown before is actually based on an iterative regression of random projections.
@@ 6.2 Numerical results
7 RRNN Security
An interesting aspect of randomness is that one can also use it to derive a simple secure scheme to train and run the RRNN model in the cloud, and later to provide access to the model. More exactly, we can encrypt (obfuscate) the data before we train the RRNN model in the cloud. The process consists in simply obfuscating the data using a secret random orthonormal matrix $Q$. Since the data is $M$ dimensional, the secret matrix $Q$ should be a random orthonormal $M\times M$ matrix. For example, such a matrix can be generated using the QR decomposition of a randomly generated matrix $\Psi$ with the elements drawn from the normal distribution [17], such that: $\Psi=QR$. The result of the decomposition is an orthonormal random matrix $Q$, and a triangular matrix $R$ which can be discarded.
@@ 7 RRNN Security
Since the data $X$ lives in a high dimensional space $\mathbb{R}^{M}$, we can encrypt (obfuscate) it before we train the RRNN (or any other neural network), as following: $E(X)=XQ$. In order to be able to use the model, the test data vector $\tilde{x}$ submitted for classification should also be encrypted using the same secret matrix: $Q\tilde{x}=\tilde{x}^{\prime}$ (here $\tilde{x}$ and $\tilde{x}^{\prime}$ are column vectors). This works because orthonormal matrices preserve the norm and the angle between data vectors. For example, all the results reported in this paper have been computed using this encryption (obfuscation) method.
Moreover, one can also derive an encrypted (obfuscated) public service by decomposing the secret matrix $Q$ as products of two orthonormal random matrices:
@@
Here the matrices $S_{n}$ and $P_{n}$ are also $M\times M$ random orthonormal matrices. The orthonormal random matrices $S_{n}$ can be generated using the same QR approach, however for each matrix $S_{n}$ one should use a different secret seed $s_{n}$ for the random number generator. The $P_{n}$ matrices are then computed using $P_{n}=S_{n}^{T}Q$, and they are also random and orthonormal, since they are the product of two random orthonormal matrices.
The host (owner) of the encrypted model can give the $P_{n}$ matrix to the user $n$, and keep the secret seed $s_{n}$ such that it can always re-generate the secret matrix $S_{n}$. Then the user $n$ can encrypt the test sample $\tilde{x}$ using $P_{n}\tilde{x}$ and send the result to the host, which also computes $S_{n}P_{n}\tilde{x}=Q\tilde{x}=\tilde{x}^{\prime}$ before it queries the model. In the next step, the host performs the classification of the encrypted $\tilde{x}^{\prime}$ using the RRNN model in the cloud, and returns the class value to the user. This way, the test sample is also encrypted during the transit, and the neural network model is rendered useless without knowing the secret matrices $S_{n}$ and $P_{n}$.
@@ 8 Conclusion
The advantage of the random-SLFNN is that the resulting training task can be formulated as a ridge regression problem. Despite their simplicity, many previous studies have indicated that such randomized neural network models can reach a sound classification performance, if they employ a large number of hidden neurons relative to the dimensionality of the samples. Here we have shown that contrary to this assumption, one can obtain good results even if the number of hidden neurons has the same order of magnitude as the dimensionality of the data samples, if the dimensionality of the samples is reasonably high. We have also exploited this idea by developing an efficient iterative training method, which significantly improves the classification accuracy. This training approach results in a deep neural network architecture, which we called Residual Random Neural Network (RRNN). The numerical results are in agreement with the theoretical results, showing that by increasing the dimensionality of the data space, the random projections drawn from the normal distribution are almost orthogonal. This surprising property of high dimensional spaces implicitly results in a much better separation of the randomly projected data samples, and consequently in an improved classification accuracy. Finally, by using orthonormal random projection matrices, we have developed an encryption (obfuscation) method which can be used to protect both the data and the neural network model. The code for RRNN is provided at: https://github.com/mandrecut/rrnn/.

### v4

@@ Abstract
The single-layer feedforward neural network with random weights is a recurring motif in the neural networks literature. The advantage of these networks is their simplified training, which reduces to solving a ridge-regression problem. A general assumption is that these networks require a large number of hidden neurons relative to the dimensionality of the data samples, in order to achieve good classification accuracy. Contrary to this assumption, here we show that one can obtain good classification results even if the number of hidden neurons has the same order of magnitude as the dimensionality of the data samples, if this dimensionality is reasonably high. Inspired by this result, we also develop an efficient iterative residual training method for such random neural networks, and we extend the algorithm to the least-squares kernel version of the neural network model. Moreover, we also describe an encryption (obfuscation) method which can be used to protect both the data and the resulted network model.
@@ 1 Introduction
Randomization of the weights in the single-layer feedforward neural network (SLFNN) is a recurring motif in the neural networks research. As a consequence, the "benefits of randomization" have been discovered (and rediscovered) many times by different authors in various contexts [1]-[9]. The main idea is the fixed random initialization of the weights in the hidden layer of the SLFNN, such that the training is reduced to computing the ridge-regression solution of the linear system corresponding to the hidden layer outputs and the classification targets. Therefore, the key advantage of these random neural networks is their simplified training algorithm. Also, it has been shown that despite of the randomness of their hidden weights layer, these simple networks are still capable of a good approximation of any non-constant piecewise continuous function [5].
A general assumption is that these random-SLFNNs require a large number of hidden neurons $J$ relative to the dimensionality of the samples $M$, in order to achieve good classification accuracy. Typically, for higher accuracy applications the number of hidden neurons is required to be an order of magnitude larger (or even more) than the dimensionality of the data samples, $J\gg M$ [1]-[9]. Thus, the classification accuracy of random-SLFNNs depends on two parameters, the dimensionality of the space $M$ and the number of the hidden random neurons $J$. Therefore, the application of random-SLFNNs to very-high-dimensional data becomes problematic, since by increasing the dimensionality of the data $M$, the number of hidden neurons $J$ is also expected to increase by an order of magnitude, or even more.
@@
Using the above intuition, here we derive a simple strategy to improve the classification accuracy of random-SLFNNs by increasing the dimensionality of the data samples $M$. This approach results in a reduced number of required hidden neurons $J$, such that both $M$ and $J$ have the same order of magnitude. We further exploit this idea by developing an efficient and low complexity iterative method, that can be used to train a residual random neural network (RRNN) model, significantly improving their classification accuracy. Following these results, we also extend the residual training algorithm to the least-squares kernel version of the neural network (RRKN), achieving a similar classification accuracy, with a more simplified architecture. Moreover, we also describe an encryption (obfuscation) method based on orthonormal random matrices, which can be used to protect both the data and the resulted network model.
@@ 4.1 Data
such that all the vectors involved in the computation lie on the $M-$dimensional unit hypersphere $\mathcal{S}^{M-1}(1)$ (an optional initial step which may improve the classification is also $x_{n}\leftarrow\sqrt{x_{n}}$ [16]). Thus, the random projection of a data sample is simply the cosine of the angle between the sample and the random projection vector. Also, the non-linear activation function used in all the numerical experiments is the hyperbolic tangent. Therefore, after accounting for the above normalization procedure, the matrices $H$ and $\tilde{H}$ from random-SLFNN equations can be rewritten as following:
@@ 5.1 Additional information
Following the conclusion of the preliminary numerical experiments, we hypothesize that by adding more information about the data samples, such that we can further increase the size of each sample, then we can expect an increase in the classification accuracy. A simple way to add more information is to compute some non-linear function of the data sample, and appended it to the initial data sample. Obviously, such a function must be non-linear, since a linear function will make the data matrix singular. For example, a simple approach is to add the absolute value of the fast Fourier transform (FFT) [16]:
@@ 5.2 Random-SLFNN with additional information
Let us now repeat the previously described experiment, by adding the additional FFT information. The results shown in Table 3 are surprisingly good, reaching an accuracy of $\eta=98.63\%$ for MNIST, and respectively $\eta=90.13\%$ for fMNIST, when $s=2$ and $q=2$. In all the considered cases the regularization parameter was set to $\lambda=0$ for MNIST, and $\lambda=10$ for fMNIST.
@@ 6.1 The algorithm
6.2 RRNN numerical results
In Figures 1 and 2 we show the numerical results obtained using the RRNN approach for the MNIST, and respectively the fMNIST datasets. One can see that for $s=2$, $q=2$ and $T^{*}=15$ iteration steps (or layers), the classification accuracy reaches $\eta\simeq 99.12\%$ for MNIST ($\lambda=0$, $\mu=1/2$), and respectively $\eta\simeq 91.41\%$ for fMNIST ($\lambda=10$, $\mu=1/2$). These results are surprisingly good for such a simple method, which as shown above is actually based on an iterative regression of random projections.
@@ 6.2 RRNN numerical results
7 Randomized kernel method
[…]
$\Phi(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle.$ (33)
An important result is that any non-linear function $f$ can be expressed as a linear combination of kernel products evaluated on the training data points $\chi=\{x_{n}|n=1,\dots,N\}$ [17]:
$f(x)=\sum_{n=1}^{N}a_{n}\Phi(x,x_{n}).$ (34)
A randomized kernel is obtained by randomly selecting a block of $J$ training samples $X_{J}\subset X$, and then computing the $N\times J$ kernel matrix [16]:
$\Phi(X,X_{J})=\varphi(\langle X,X_{J}\rangle),$ (35)
where $\varphi$ is an element wise nonlinear function of the inner product:
$\langle X,X_{J}\rangle=XX_{J}^{T}.$ (36)
The weights of the output layer are determined by solving the least-squares regression problem:
$W_{J}=\text{arg}\min_{W_{J}}\left\{\|\Phi(X,X_{J})W_{J}-Y\|^{2}+\lambda\|W_{J}\|^{2}\right\},$ (37)
such that, the weights $W_{J}$ are:
$W_{J}=\left[\Phi(X,X_{J})\Phi(X,X_{J})^{T}+\lambda I\right]^{-1}\Phi(X,X_{J})^{T}Y.$ (38)
Here, $I$ is the $J\times J$ identity matrix, $\lambda\geq 0$ is the regression parameter, and each row $y_{n}\in\mathbb{R}^{K}$ of $Y$ corresponds to the class of the data point $x_{n}$.
In the case of RRKN the classes are encoded using the same one-hot encoding procedure (4)-(5), however in this case the target approximations are given by:
$\tilde{Y}=\Phi(\tilde{X},X_{J})^{T}W_{J}=[\tilde{y}_{i,k}]_{N^{\prime}\times K}.$ (39)
We should emphasize that in the classification step, $\Phi(\tilde{X},X_{J})$ is the kernel matrix computed using the test data $\tilde{X}$ and the same randomly selected block of training data $X_{J}$.
8 RRKN approach
8.1 The algorithm
As in the case of RRNNs, the main mechanism of the residual approach is to iteratively subtract the predicted values $\hat{Y}$ from the target values $Y$, and then to re-apply the method again on the resulted residual:
  1. 1.
    Set the maximum number of iterations $T$, the residual limit $\varepsilon>0$, the number $J$ of the samples in the random kernel, the regression parameter $\lambda\geq 0$, and the seed of the random number generator.
  2. 2.
    Set the iteration index to $t=0$, and the initial residual $R(0)=Y$.
  3. 3.
    Randomly select the block matrix $X_{J}(t)$ from $X$.
  4. 4.
    Compute the kernel matrix: $\Phi(X,X_{J}(t))$. Here we use a polynomial kernel with $\Phi(X,X_{J}(t))=\langle X,X_{J}(t)\rangle^{(k)}$, where $(k)$ means element-wise power. Best results were obtained for $k=5$.
  5. 5.
    Solve the regression problem:
    $W_{J}(t)=\left[\Phi(X,X_{J}(t))\Phi(X,X_{J}(t))^{T}+\lambda I\right]^{-1}\Phi(X,X_{J}(t))^{T}R(t)$
  6. 6.
    Compute the residual: $R(t+1)=R(t)-\Phi(X,X_{J}(t))W_{t}$
  7. 7.
    If $\|R(t+1)\|<\varepsilon$ or $t+1>T$ then set $T^{*}\leftarrow t$ and stop,
    otherwise set $t\leftarrow t+1$ and go to 3.
  8. 8.
    Return $W_{J}(t)$ for $t=0,1,...,T^{*}$.
One can see that at each time step $t=0,1,...,T^{*}$ we compute a new kernel matrix using a randomly selected projection matrix $X_{J}(t)$, and a new residual $R(t+1)$. Finally, the weight matrices $W_{J}(t)$ can be used to find the target approximation using the $\tilde{X}$ test data matrix:
$\tilde{Y}=\sum_{t=0}^{T^{*}}\Phi(\tilde{X},X_{J}(t))W_{t}.$ (40)
We notice again that since the kernel projection matrices $X_{J}(t)$ are randomly selected, we do not have to store them, we only need the random seed. However, as in the case of RRNNs we do need to store the weights $W_{J}(t)$. Thus, if $T^{*}$ is the number of steps at the end of the iteration, then the training process will produce a RRKN with $T^{*}$ corrective layers.
8.2 RRKN numerical results
In the first experiment we scale the kernel matrices as following $J=s1000$, $s=1,2,5,10,15$. For each $J$ value, we limit the number of approximation steps (layers) to $T=20$, and we let be their index $t\in\{1,2,...,20\}$. The results for the MNIST data are shown in Figure 3, and respectively for the fMNIST data in Figure 4. One can see that the classification accuracy improves gradually by increasing the number of approximation layers $T$, and the size $J$ of the kernel matrix, reaching the best values for $T=20$ and $J=15000$: $98.87\%$ for MNIST, and respectively $91.04\%$ for fMNIST.
Figure 3: Initial RRKN results for MNIST: (a) accuracy; (b) residual norm.
Figure 4: Initial RRKN results for fMNIST: (a) accuracy; (b) residual norm.
In the second experiment we increase the size of the images by a factor of two, using the same "resize" method from the "scikit-image" package in Python. The results for the MNIST data are shown in Figure 5, and respectively for the fMNIST data in Figure 6. The classification accuracy again improves, and in this case the best results are $99.04\%$ for MNIST, and respectively $91.34\%$ for fMNIST.
Figure 5: Image resizing RRKN results for MNIST: (a) accuracy; (b) residual norm.
Figure 6: Image resizing RRKN results for fMNIST: (a) accuracy; (b) residual norm.
As mentioned in the case of RRNNs, instead of simply resizing the images, a better approach is to append more information to them. Therefore, we add again the absolute value of the fast Fourier transform (FFT) to each image (30)-(31), resulting in a row data vector $(x_{n},x^{\prime}_{n})$, with a length of $23M/$ 2. We should emphasize that in this case the length of the data vectors is $23M/$ 2, which is much smaller than length resulted in the second experiment (Figures 5 and 6), where we basically quadrupled the length of the images, to $4M$, as a consequence of the scaling procedure. Obviously, adding new information to each data sample (in order to increase dimensionality) has a more significant effect than simply resizing the images.
In this case, the numerical results for the same parameters are shown in Figure 7 for MNIST, and respectively in Figure 8 for fMNIST. One can see that the best results are $99.11\%$ for MNIST and $91.43\%$ for fMNIST, which are comparable to the results previously reported for the RRNNs. Also, one can see that in all cases, the classification accuracy of the residual method is starting to saturate for $J>10000$ and $t>15$.
Figure 7: RRKN results for MNIST with additional FFT data: (a) accuracy; (b) residual norm.
Figure 8: RRKN results for fMNIST with additional FFT data: (a) accuracy; (b) residual norm.
9 Security approach
@@
An interesting aspect of randomness is that one can also use it to derive a simple secure scheme to train and run the RRNN/RRKN model in the cloud, and later to provide access to the model. More exactly, we can encrypt (obfuscate) the data before we train the models. The process consists in simply obfuscating the data using a secret random orthonormal matrix $Q$. Since the data is $M$ dimensional, the secret matrix $Q$ should be a random orthonormal $M\times M$ matrix. For example, such a matrix can be generated using the QR decomposition of a randomly generated matrix $\Psi$ with the elements drawn from the normal distribution [18], such that: $\Psi=QR$. The result of the decomposition is an orthonormal random matrix $Q$, and a triangular matrix $R$ which can be discarded. Since the data $X$ lives in a high dimensional space $\mathbb{R}^{M}$, we can encrypt (obfuscate) it before we train the RRNN/RRKN model (or any other neural network), as following: $E(X)=XQ$. In order to be able to use the model, the test data vector $\tilde{x}$ submitted for classification should also be encrypted using the same secret matrix: $Q\tilde{x}=\tilde{x}^{\prime}$ (here $\tilde{x}$ and $\tilde{x}^{\prime}$ are column vectors). This works because orthonormal matrices preserve the norm and the angle between data vectors. For example, all the results reported in this paper have been computed using this encryption (obfuscation) method. Moreover, one can also derive an encrypted (obfuscated) public service by decomposing the secret matrix $Q$ as products of two orthonormal random matrices:
@@
The matrices $S_{n}$ and $P_{n}$ are also $M\times M$ random orthonormal matrices. The orthonormal random matrices $S_{n}$ can be generated using the same QR approach, however for each matrix $S_{n}$ one should use a different secret seed $s_{n}$ for the random number generator. The $P_{n}$ matrices are then computed using $P_{n}=S_{n}^{T}Q$, and they are also random and orthonormal, since they are the product of two random orthonormal matrices.
The host (owner) of the encrypted model can give the $P_{n}$ matrix to the user $n$, and keep the secret seed $s_{n}$ such that it can always re-generate the secret matrix $S_{n}$. Then the user $n$ can encrypt the test sample $\tilde{x}$ using $P_{n}\tilde{x}$ and send the result to the host, which also computes $S_{n}P_{n}\tilde{x}=Q\tilde{x}=\tilde{x}^{\prime}$ before it queries the model. In the next step, the host performs the classification of the encrypted $\tilde{x}^{\prime}$ using the RRNN/RRKN model in the cloud, and returns the class value to the user. This way, the test sample is also encrypted during the transit, and the neural network model is rendered useless without knowing the secret matrices $S_{n}$ and $P_{n}$.
@@ 10 Conclusion
The advantage of the random-SLFNN is that the resulting training task can be formulated as a ridge regression problem. Despite their simplicity, many previous studies have indicated that such randomized neural network models can reach a sound classification performance, if they employ a large number of hidden neurons relative to the dimensionality of the samples. Here we have shown that contrary to this assumption, one can obtain good results even if the number of hidden neurons has the same order of magnitude as the dimensionality of the data samples, if the dimensionality of the samples is reasonably high. We have also exploited this idea by developing an efficient iterative training method, which significantly improves the classification accuracy. This training approach results in an error correcting neural network architecture, which we called Residual Random Neural Network (RRNN). The numerical results are in good agreement with the theoretical results, showing that by increasing the dimensionality of the data space, the random projections drawn from the normal distribution are almost orthogonal. This surprising property of high dimensional spaces implicitly results in a much better separation of the randomly projected data samples, and consequently in an improved classification accuracy. Following these results we have also extended the residual training algorithm to the least-squares Residual Random Kernel Network (RRKN), achieving a similar classification accuracy performance, with a more simplified architecture. Finally, by using orthonormal random projection matrices, we have developed an encryption (obfuscation) method which can be used to protect both the data and the resulted network model. The code for the RRNN/RRKN models is provided at: https://github.com/mandrecut/rrnn/.